OpenStudy (firejay5):

x = 7y + 2 in standard form

OpenStudy (anonymous):

:-)

OpenStudy (firejay5):

:(

OpenStudy (geometry_hater):

Know the formula for Standard form?

OpenStudy (anonymous):

I don't either, tell me...

OpenStudy (firejay5):

yes I do

OpenStudy (geometry_hater):

lol you and your sarcasm!!! Its A squared + B squared = c squared

OpenStudy (firejay5):

take off the squared from C

OpenStudy (anonymous):

If you said "linear equation" or "line" that might help...

OpenStudy (anonymous):

get the x and y on the same side of the equal sign and clear fractions. Formula is Ax +By =C

OpenStudy (firejay5):

not helping me

OpenStudy (anonymous):

how can I help? do this first subtract 7y from both sides

OpenStudy (anonymous):

getting x-7y=2

OpenStudy (anonymous):

In linear algebra, it is customary to write Ax + By = C as the geometry is secondary. For a line, I prefer f(x) = y = mx + c

OpenStudy (anonymous):

now ask your self is it in this form Ax+By=C

OpenStudy (firejay5):

it's not

OpenStudy (anonymous):

why?

OpenStudy (firejay5):

cause it's not

OpenStudy (anonymous):

it is, look at this

OpenStudy (firejay5):

I need it in standard form and it's not

OpenStudy (anonymous):



OpenStudy (firejay5):

How did you get that I am asking

OpenStudy (anonymous):

standard form for what? there are many standard forms for

OpenStudy (anonymous):

you started with this right x=7y+2

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

I did and...

OpenStudy (anonymous):

x = 7y + 2 Subtract 7y -> x - 7y = 2 is the so-called standard form.

OpenStudy (firejay5):

yes that is correct

OpenStudy (firejay5):

estudier

OpenStudy (anonymous):

I think its bs personally......:-)

OpenStudy (firejay5):

Did you learn it differently???

OpenStudy (firejay5):

why do you think it's bull pellet

OpenStudy (anonymous):

When I learned (a long time ago) I just learned all the different possible representations for a line. I think it depends on whatever you are doing, maybe sometimes its convenient to have a vector equation or a parametric equation.

OpenStudy (firejay5):

what's you real name (first+last name)

OpenStudy (anonymous):

Textbooks dont call this standard form anymore

OpenStudy (anonymous):

Good!

OpenStudy (firejay5):

what's your name estudier

OpenStudy (anonymous):

If I wanted people to know that, I would have registered with it.....:-)

OpenStudy (firejay5):

do you have facebook, because you are really good at this

OpenStudy (anonymous):

I have but I only use for site logins.....

OpenStudy (firejay5):

could I have it so we could stay in touch